OpenStudy (anonymous):

Please help with give medal. You begin solving the equation 4 – 5x = 59 by subtracting 4 from both sides. Which is the best choice for Step 2? A) Subtract 55 from both sides B) Multiply both sides by -5 C) Divide both sides by 55 D) Divide both sides by -5

OpenStudy (texaschic101):

4 - 5x = 59 --- subtract 4 from both sides -5x = 59 - 4 -5x = 55 -- divide both sides by -5 x = -11

OpenStudy (anonymous):

Thank you that makes sense now.

OpenStudy (anwaarullah):

D